# TalkingEyes: Pluralistic Speech-Driven 3D Eye Gaze Animation

### v1

@@ III-B1 Video pre-processing
The visibility of the subject’s eyes in video is paramount for the subsequent 3D eye gaze fitting. Hence, the videos need be excluded if the eyes are obscured. For videos where the invisibility of eyes is caused by head turning, we perform a preliminary exclusion by detecting significant disparities in the distance between the left and right eye corners and the side of the face. Then, the videos featuring individuals with obstructions such as sunglasses, which inherently mask the eyes, are excluded by manual screening.
@@
We perform frame-by-frame reconstruction on videos to generate the 3D facial motion and 3D head motion. In order to obtain the highest quality possible, we utilize the 3D face reconstruction network proposed in EMOTE [23], which combines the strengths of four SOTA methods: EMOCA[54], SPECTRE[55], DECA[56] and MICA[57]. All the four SOTA methods use FLAME as 3D face representation. On the output side, DECA outputs the global rotation of head, forming the 3D head motion; MICA outputs the facial shape vector, EMOCA outputs the expression vector and DECA outputs the jaw pose, together forming the 3D facial motion. SPECTRE is only used in network training to enhance lip articulation by lip-reading loss.
@@ III-B4 Post-processing
Post-processing is needed to refine the motion data obtained in the previous step. When processing frames that contain blinking, we maintain the eye pose based on the last known open-eye state. This is because the iris’s position cannot be reliably detected during a blink. By doing so, we can ensure the continuity and naturalness of the facial animation, effectively preventing sudden movement of eyes that might occur during a blink event. Then, we conduct temporal convolutional smoothing individually on the three types of motion data to reduce unnatural jitter and combine them in one FLAME model.
@@ III-C LightGazeFit
The deep-learning-based gaze estimation methods [43] proposed in computer vision provide a number of potential solutions for reconstructing eye gaze from video, but their estimated 3D gaze directions are not always compatible with the reconstructed 3D face mesh. Alternatively, 3D eye gaze fitting methods [40, 41, 42] from computer graphics seem to offer a better fit, but their code has not been publicly released.
@@
The proposed LightGazeFit is a lightweight yet effective 3D eye gaze fitting method tailored for processing videos from VoxCeleb2 and HDTF. Existing fitting methods [40, 41] heavily rely on accurate segmentation of iris and pupil pixels to formulate mask or edge likelihood. However, due to the relatively poor quality of VoxCeleb2 videos, which are only available in a $224\times 224$ resolution, achieving precise segmentation of iris and pupil is nearly impossible. LightGazeFit overcomes this issue by utilizing only the pupil center to form the likelihood, yet it still generates good results across all videos from VoxCeleb2 and HDTF.
The 3D eyeball model used in LightGazeFit comes from FLAME and it’s shown in Fig. 4. We manually annotate the iris vertices, the pupil vertices and the pupil center vertex on the 3D model. Then, we assign a brown color to the iris vertices and a black color to the pupil vertices. As the precise segmentation of iris and pupil is not available, LightGazeFit eliminates the need of eyeball calibration [40, 41, 42] which estimates the iris and pupil size for each subject. With the iris and pupil size fixed, we found that it performs well in the eyeball rotation estimation.
@@
where $\lambda_{1}$ and $\lambda_{2}$ are hyper-parameters and are set to $41e-$ 4 and $21e-$ 2 respectively. The pupil center loss ${\mathcal{L}_{\text{center}}}$ is defined as the difference between the predicted 2D pupil center $\mathbf{y}$ and the point $\hat{\mathbf{y}}$ projected from the pupil center vertex on 3D eyeball model to the 2D image plane:
$\mathcal{L}_{\text{center}}=||\hat{\mathbf{y}}-\mathbf{y}||_{2}^{2}.$ (2)
The 2D pupil center $\mathbf{y}$ in the image is predicted by a neural network [41]. Given the 2D pupil center $\mathbf{y}$, their exists many 3D eyeball rotation solutions to generate $\hat{\mathbf{y}}$. To reduce such ambiguity, we use regularization loss to make the eyeball rotate as little as possible. The regularization loss $\mathcal{L}_{\text{reg}}$ is defined as:
$\mathcal{L}_{\text{reg}}=||\mathbf{p}_{eye}||_{2}^{2},$ (3)
@@
$\mathcal{L}_{\text{sym}}=||\mathbf{p}_{eye\_l}-\mathbf{p}_{eye\_r}||_{2}^{2}.$ (4)
@@ IV Method: TalkingEyes
Our ultimate goal is to synthesize eye gaze motion, eye blinks, head motion and facial motion collectively from the speech, under the unified 3D face representation FLAME [34]. For facial motion, we directly use Learn2Talk [19] as the generator which has achieved impressive 3D lip-sync performance. The eye gaze motion and head motion are jointly generated by a cascade VAE-VQVAE mechanism and a temporal autoregressive model. We introduce the motivation of using VAE and VQVAE in Sect. IV-A, the pipeline of TalkingEyes in Sect. IV-B, the two main stages in training in Sect. IV-C and IV-D, and the synthesis of eye blinks in Sect. IV-E.
@@
IV-A Latent Spaces Learning by VAE-VQVAE
To model the weak correlation between speech signals and non-verbal cues in conversation, such as head gesture [58], hand gesture [59, 60] and body poses [61], the generative models VAE [1, 62] and VQVAE [2, 63] are commonly used. These models are capable of learning a compact representation for the non-verbal motion data. Through mapping the speech to the compact latent space, the difficulty in capturing the weak correlation between speech and motion is significantly attenuated and hence promotes the quality of motion synthesis. Following this, TalkingEyes also employ VAE and VQVAE to learn the latent space of motion.
On the other hand, there is a key difference between VAE and VQVAE in terms of motion diversity. VAE tends to output the mean of a Gaussian distribution, thus limiting the diversity in outputs. VQVAE uses discrete codebook to capture a diverse set of latent representations, allowing for a wider range of possible outputs. Since the rotation range of head usually exceeds that of eyeballs, the variation of the head motions is generally higher than that of eye gaze motions. As summarized in Tab. II, VAE is capable of producing head motions with adequate diversity (e.g. PoseVAE [58, 19]), while VQVAE may sometimes produce excessive diverse and exaggerated head motions. In contrast, VAE struggles to produce sufficiently diverse eye gaze motions, while VQVAE effectively achieve this. To support our claims, we have conducted quantitative experiments which are introduced in Sect. V-D. To conclude, we use VAE in the latent space learning for head motions, and use VQVAE in that for eye gaze motions.
@@ IV-B Pipeline
The training of TalkingEyes comprises two main stages: firstly learn a discrete latent space for eye gaze motions by pre-training VQVAE, and then jointly train the VAE-based speech-to-head translation and the VQVAE-based speech-to-gaze translation in an autoregressive manner. In the first stage, we pre-train a VQVAE to model the latent space of eye gaze motions as a discrete codebook $\mathcal{Z}$. The codebook is learned by self-reconstruction over the ground truth motions $\mathbf{P}^{e}_{1:T}$. In the second stage, a conditional VAE is employed to reconstruct the head motions $\mathbf{\hat{P}}^{h}_{1:T}$ from the ground truth motions $\mathbf{P}^{h}_{1:T}$. This motion reconstruction is conditioned on the speech embedding $\mathbf{A}_{1:T}$ that is extracted by the speech encoder (wav2vec2.0 [64]) from the input speech. Then the predicted head motions $\mathbf{\hat{P}}^{h}_{1:T}$ and the previously predicted eye gaze motions $\mathbf{\hat{P}}^{e}_{1:t}$ are used as conditions in the mapping of the speech embedding $\mathbf{A}_{1:T}$ to the target motion codes. This mapping is accomplished through multi-modal alignment and codebook $\mathcal{Z}$ lookup. The motion codes are further decoded into the eye gaze motion $\mathbf{\hat{P}}^{e}_{t+1}$ at current frame $t+1$, which will serve as one frame of past motions in the next round of autoregression.
@@ IV-C Discrete Latent Space of Eye Gaze Motion
Similar to [2, 16], the training losses of VQVAE include a motion-level loss and two intermediate code-level losses:
@@ IV-D Speech-to-head and Speech-to-gaze Translations
$\mathcal{L}_{\text{rec}}=||\hat{\mathbf{P}}^{h}_{1:T}-\mathbf{P}^{h}_{1:T}||_{2}^{2}+||\hat{\mathbf{P}}^{e}_{1:T}-\mathbf{P}^{e}_{1:T}||_{2}^{2}.$ (11)
@@
$\begin{split}\mathcal{L}_{\text{vel}}=\sum_{t=1}^{T-1}||(\mathbf{\hat{P}}^{h}_{t+1}-\mathbf{\hat{P}}^{h}_{t})-(\mathbf{P}^{h}_{t+1}-\mathbf{P}^{h}_{t})||_{2}^{2}\\ +\sum_{t=1}^{T-1}||(\mathbf{\hat{P}}^{e}_{t+1}-\mathbf{\hat{P}}^{e}_{t})-(\mathbf{P}^{e}_{t+1}-\mathbf{P}^{e}_{t})||_{2}^{2}.\end{split}$ (12)
The KL-divergence loss ${\mathcal{L}_{\text{kl}}}$ forces the Gaussian distribution of the predicted head motions $\mathbf{P}^{h}_{1:T}\sim\mathcal{N}(\mu_{h},\sigma_{h}^{2})$ close to a standard Gaussian $\mathcal{N}(0,1)$. ${\mathcal{L}_{\text{kl}}}$ has a simplified form [1]:
@@ IV-E Eye Blink
To generate more vivid animation, we have also implemented the synthesis of eye blinks. Blinking is primarily influenced by personal habits, environmental factors and other conditions. Liu et al. [65] counted the number of blinks in their dataset and found that the frequency of human blinking, measured in blinks per minute, typically follows a log-normal distribution. Following their work, we conducted a statistical analysis on our dataset TKED to derive a suitable blink frequency that well aligns with our eye gaze animation. We counted the number of blinks in each video in TKED and then converted these counts into blinks per minute based on the video’s duration. For blink detection, we calculated the Eye Aspect Ratio (EAR) [66], which is defined as the ratio of the eye’s height to its width. When the eyes are closed, the EAR approaches zero. After statistical analysis, we found that the blinks per minute follow a Gaussian distribution with a mean of 40.10 and a deviation of 15.42. When generating animations, we first sample the number of blinks per minute from the distribution and then convert that into blink intervals. Subsequently, we adjust the expression parameters within a 5-frame window for each blink to create the blinking action.
@@ V-A Training Details
VQVAE is pre-trained by Adam optimizer for 100 epochs with a learning rate $41e-$ 4. The pre-training takes about $1$ hour on a NVIDIA RTX 4090 GPU. Then the VAE and the autoregressive model are trained by Adam optimizer for 100 epochs with a learning rate $41e-$ 4. The training takes about $5$ hours on the same GPU.
@@ V-B2 Baselines
Since the recent speech-driven 3D eye gaze animation methods are scarce and early computer animation methods [29, 30, 31, 32, 33] targeted at this task have not released their codes or datasets, we implemented four 3D eye gaze motion generation methods as the baselines for quantitative comparison.
@@ V-B3 Evaluation Metrics
Figure 6: Visual comparisons of 3D eye gaze motions generated from input audio by using “Mean”, “Random”, “Similar Sample”, “VAE” and our method.
V-C Evaluation of TalkingEyes
V-C1 Quantitative Comparison
Tab. III shows the quantitative evaluation results over TKED-Test. Our TalkingEyes outperforms other methods in terms of motion diversity and audio-motion correlation. TalkingEyes is capable of generating more diverse motions than the “Random” selection, thanks to VQVAE’s ability to learn a structured and clustered representation from the original data. Additionally, TalkingEyes achieves the highest score in audio-motion correlation, indicating that mapping the audio signal to the discrete motion space is a more effective and robust way for modeling the weak relationship between audio and eye gaze motion compared to other approaches, such as searching for similar samples in the audio feature space. When it comes to $L_{2}$ distance, TalkingEyes slightly compromises accuracy in matching the ground truth, which is a common trade-off in non-deterministic tasks [67, 68]. It is worth noting that “VAE” achieves very similar performance to that of “Mean” with respect to the three evaluation metrics. This further supports our motivation in the choice of VAE or VQVAE, as discussed in Sect. IV-A: VAE tends to output the mean of the Gaussian constructed on the eye gaze motions.
Ours vs. Baseline Realism
Ours vs. Mean 85 (3.3e-9)
Ours vs. Random 86 (7.5e-15)
Ours vs. Similar Sample 93 (5.7e-22)
Ours vs. VAE 83 (5.7e-10)
TABLE IV: User study results. The percentage of answers where our method is preferred over the competitor is listed. The numbers in the brackets denote p-value.
@@ V-C2 Perceptual User Study
For more comprehensive evaluation, we have conducted a perceptual user study to evaluate the realism of the animated 3D eye gaze. Specifically, we adopt the A/B test in a random order to compare our TalkingEyes with the four baselines. In each comparison, the participants were asked to answer the question with A or B: which recording exhibits more natural eyeball movements with fewer artifacts, such as exaggerated movements, static poses and motion jitter? We randomly selected 20 audios from TKED-Test, and used them to generate animations in the four kinds of comparisons (5 audios $\times$ 4 comparisons). 20 participants took part in the study, finally yielding 400 entries. We calculated the ratio of participants who prefer our method over the baseline. The percentage of A/B testing is tabulated in Tab. IV, which shows that participants favor our method over all the four baselines. We have also conducted ANOVA tests in the user study, in which the p-value less than 0.05 indicates a statistically significant difference between our method and the competitor.
Figure 7: Visual comparisons of sampled frames of 3D facial animations generated by FaceFormer [15], CodeTalker [16], Learn2Talk [19] and our method.

### v2

@@ III-B1 Video Pre-processing
Clear visibility of the subject’s eyes in video is paramount for the subsequent 3D eye gaze fitting. Hence, we excluded videos where the eyes are obscured. For instances of poor visibility due to head turns, we performed a preliminary exclusion by detecting significant disparity in the distances from the left and right eye corners to the two sides of the face. Additionally, the videos featuring individuals with obstructions such as sunglasses, which inherently mask the eyes, were excluded by manual screening. After this pre-processing step, the number of the raw video clips was reduced from approximately 20,000 to 6,000.
@@
We performed frame-by-frame reconstruction on videos to generate the 3D facial motion and 3D head motion. In order to obtain the highest quality possible, we utilized the 3D face reconstruction network proposed in EMOTE [EMOTE2023], which combines the strengths of four SOTA methods: EMOCA[EMOCA2022], SPECTRE[LipreadVideo2022], DECA[DECA2021] and MICA[zielonka2022towards]. All the four SOTA methods use FLAME as 3D face representation. On the output side, DECA outputs the global rotation of head, forming the 3D head motion; MICA outputs the facial shape vector, EMOCA outputs the expression vector and DECA outputs the jaw pose, together forming the 3D facial motion. SPECTRE is only used in network training to enhance lip articulation by lip-reading loss.
@@ III-B4 Post-processing
Post-processing refines the motion data obtained from the previous steps. For frames containing blinks, the eye pose was maintained from the last known open-eye state because the iris position can’t be reliably detected during a blink. This ensures the continuity and naturalness of the facial animation by preventing unnatural eye movements during blinks. We then applied temporal convolutional smoothing individually to the three types of motion data to reduce jitter and finally combined them in a single FLAME model.
@@ III-C LightGazeFit
The deep-learning-based gaze estimation methods [GazeEstimationSurvey2024] proposed in computer vision provide a number of potential solutions for reconstructing eye gaze from video, but their estimated 3D gaze directions are not always compatible with the reconstructed 3D face mesh. Alternatively, 3D eye gaze fitting methods [wang2016realtime, wang2019realtime, wen2017real] from computer graphics seem to offer a better fit, but their code has not been publicly released.
@@
The proposed LightGazeFit is a lightweight yet effective 3D eye gaze fitting method tailored for processing videos from VoxCeleb2 and HDTF. Existing fitting methods [wang2016realtime, wang2019realtime] heavily rely on accurate segmentation of iris and pupil pixels to formulate mask or edge likelihood. However, due to the relatively poor quality of VoxCeleb2 videos, which are only available in a $224\times 224$ resolution, achieving precise segmentation of iris and pupil is nearly impossible. LightGazeFit overcomes this issue by utilizing only the pupil center to form the likelihood, yet it still generates good results across all videos from VoxCeleb2 and HDTF.
The 3D eyeball model used in LightGazeFit comes from FLAME and it’s shown in Fig. 4. We manually annotated the iris vertices, the pupil vertices and the pupil center vertex on the 3D model. Then, we assigned a brown color to the iris vertices and a black color to the pupil vertices. As the precise segmentation of iris and pupil is not available, LightGazeFit eliminates the need of eyeball calibration [wang2016realtime, wang2019realtime, wen2017real] which estimates the iris and pupil size for each subject. With the iris and pupil size fixed, we found that it performs well in the eyeball rotation estimation.
@@
where $\lambda_{1}$ and $\lambda_{2}$ are hyper-parameters and are set to 1e41e-4 and 1e21e-2 respectively. The pupil center loss $\mathcal{L}_{\text{center}}$} is defined as the difference between the predicted 2D pupil center $\mathbf{y}$ and the point $\hat{\mathbf{y}}$ projected from the pupil center vertex on 3D eyeball model to the 2D image plane:
[…]
The 2D pupil center $\mathbf{y}$ in the image was predicted by a neural network [wang2019realtime]. Given the 2D pupil center $\mathbf{y}$, their exists many 3D eyeball rotation solutions to generate $\hat{\mathbf{y}}$. To reduce such ambiguity, we used regularization loss to make the eyeball rotate as little as possible. The regularization loss $\mathcal{L}_{\text{reg}}$ is defined as:
[…]
@@
[…]
We used the Adam optimizer with a learning rate of 5e35e-3 to solve this non-linear least square optimization problem defined as Eq. 1.
@@ IV Method: TalkingEyes
Our ultimate goal is to collectively synthesize eye gaze motion, eye blinks, head motion and facial motion from speech, under the unified 3D face representation FLAME [FLAME2017]. For facial motion, we directly use Learn2Talk [Learn2Talk2025] as the generator due to its impressive 3D lip-sync performance. Eye gaze motion and head motion are firstly transformed into their respective latent spaces via VQVAE and VAE, and are then jointly generated by a temporal autoregressive model. We detail the motivation of using VAE and VQVAE in Sect. IV-A, the pipeline of TalkingEyes in Sect. IV-B, the two main stages in training in Sect. IV-C and IV-D, and the synthesis of eye blinks in Sect. IV-E.
@@
IV-A Latent Spaces Learning by VAE and VQVAE
To model the weak correlation between speech signals and non-verbal cues in conversation, such as head gesture [SadTalker2023], hand gesture [TalkShow2023, GestureDiffuCLIP2023] and body poses [EMAGE2024], the generative models VAE [vae2014, vae2016] and VQVAE [VQVAE2017, VQVAE2-2019] are commonly used. These models are capable of learning a compact representation for the non-verbal motion data. Through mapping the speech to the compact latent space, the difficulty in capturing the weak correlation between speech and motion is significantly attenuated and hence promotes the quality of motion synthesis. Following this, TalkingEyes also employ VAE and VQVAE to learn the latent space of motion.
On the other hand, there is a key difference between VAE and VQVAE in terms of motion diversity. VAE tends to output the mean of a Gaussian distribution, thus limiting the diversity in outputs. VQVAE uses discrete codebook to capture a diverse set of latent representations, allowing for a wider range of possible outputs. Since the rotation range of head usually exceeds that of eyeballs, the variation of the head motions is generally higher than that of eye gaze motions. As summarized in Tab. II, VAE is capable of producing head motions with adequate diversity (e.g. PoseVAE [SadTalker2023, Learn2Talk2025]), while VQVAE may sometimes produce excessive diverse and exaggerated head motions. In contrast, VAE struggles to produce sufficiently diverse eye gaze motions, while VQVAE effectively achieve this. To support our claims, we have conducted quantitative experiments which are introduced in Sect. V-D. To conclude, we use VAE in the latent space learning for head motions, and use VQVAE in that for eye gaze motions.
@@ IV-B Pipeline
The training of TalkingEyes comprises two main stages: firstly learn a discrete latent space for eye gaze motions by pre-training VQVAE, and then jointly train the VAE-based speech-to-head translation and the VQVAE-based speech-to-gaze translation in an autoregressive manner. In the first stage, we pre-train a VQVAE to model the latent space of eye gaze motions as a discrete codebook $\mathcal{Z}$. The codebook is learned by self-reconstruction over the ground truth motions $\mathbf{P}^{e}_{1:T}$. In the second stage, a conditional VAE is employed to reconstruct the head motions $\mathbf{\hat{P}}^{h}_{1:T}$ from the ground truth motions $\mathbf{P}^{h}_{1:T}$. This motion reconstruction is conditioned on the speech embedding $\mathbf{A}_{1:T}$ that is extracted by the speech encoder (wav2vec2.0 [wav2vec2020]) from the input speech. Then the predicted head motions $\mathbf{\hat{P}}^{h}_{1:T}$ and the previously predicted eye gaze motions $\mathbf{\hat{P}}^{e}_{1:t}$ are used as conditions in the mapping of the speech embedding $\mathbf{A}_{1:T}$ to the target motion codes. This mapping is accomplished through multi-modal alignment and codebook $\mathcal{Z}$ lookup. The motion codes are further decoded into the eye gaze motion $\mathbf{\hat{P}}^{e}_{t+1}$ at current frame $t+1$, which will serve as one frame of past motions in the next round of autoregression.
@@ IV-C Discrete Latent Space of Eye Gaze Motion
Similar to [VQVAE2017, CodeTalker2023], the training losses of VQVAE include a motion-level loss and two intermediate code-level losses:
@@ IV-D Speech-to-head and Speech-to-gaze Translations
[…]
@@
[…]
The KL-divergence loss $\mathcal{L}_{\text{kl}}$} forces the Gaussian distribution of the predicted head motions $\mathbf{P}^{h}_{1:T}\sim\mathcal{N}(\mu_{h},\sigma_{h}^{2})$ close to a standard Gaussian $\mathcal{N}(0,1)$. $\mathcal{L}_{\text{kl}}$} has a simplified form [vae2014]:
@@ IV-E Eye Blink
To generate more vivid animation, we have also implemented the synthesis of eye blinks. Blinking is primarily influenced by personal habits, environmental factors and other conditions. Liu et al. [liu2024emoface] counted the number of blinks in their dataset and found that the frequency of human blinking, measured in blinks per minute, typically follows a log-normal distribution. Following their work, we conducted a statistical analysis on our dataset TKED to derive a suitable blink frequency that well aligns with our eye gaze animation. We counted the number of blinks in each video in TKED and then converted these counts into blinks per minute based on the video’s duration. For blink detection, we calculated the Eye Aspect Ratio (EAR) [cech2016real], which is defined as the ratio of the eye’s height to its width. When the eyes are closed, the EAR approaches zero. After statistical analysis, we found that the blinks per minute follow a Gaussian distribution with a mean of 40.10 and a deviation of 15.42. When generating animations, we first sample the number of blinks per minute from the distribution and then convert that into blink intervals. Subsequently, we adjust the expression parameters within a 5-frame window for each blink to create the blinking action.
@@ V-A Training Details
VQVAE is pre-trained by Adam optimizer for 100 epochs with a learning rate 1e41e-4. The pre-training takes about $1$ hour on a NVIDIA RTX 4090 GPU. Then the VAE and the autoregressive model are trained by Adam optimizer for 100 epochs with a learning rate 1e41e-4. The batch size is set to 1. The training takes about $5$ hours on the same GPU.
@@ V-B2 Baselines
Since the recent speech-driven 3D eye gaze animation methods are scarce and early computer animation methods [HeadEyeAnimation2007, VirtualCharacterSCA2013, ConversationalAgents2012, LiveSpeehEyeTVCG2012, EyeMotion2019] targeted at this task have not released their codes or datasets, we implemented four 3D eye gaze motion generation methods as the baselines for quantitative comparison.
@@ V-B3 Evaluation Metrics
V-C Evaluation of TalkingEyes
V-C1 Quantitative Comparison
Tab. III shows the quantitative evaluation results over TKED-Test. Our TalkingEyes outperforms other methods in terms of motion diversity and audio-motion correlation. TalkingEyes is capable of generating more diverse motions than the “Random” selection, thanks to VQVAE’s ability to learn a structured and clustered representation from the original data. Additionally, TalkingEyes achieves the highest score in audio-motion correlation, indicating that mapping the audio signal to the discrete motion space is a more effective and robust way for modeling the weak relationship between audio and eye gaze motion compared to other approaches, such as searching for similar samples in the audio feature space. When it comes to $L_{2}$ distance, TalkingEyes slightly compromises accuracy in matching the ground truth, which is a common trade-off in non-deterministic tasks [PICCVPR2019, ICTPAMI2024]. It is worth noting that “VAE” achieves very similar performance to that of “Mean” with respect to the three evaluation metrics. This further supports our motivation in the choice of VAE or VQVAE, as discussed in Sect. IV-A: VAE tends to output the mean of the Gaussian constructed on the eye gaze motions.
@@ V-C1 Quantitative Comparison
Method A vs. Method B Preference Rate
Ours vs. Mean 81 (1.68e-4)
Ours vs. Random 80 (1.34e-4)
Ours vs. Similar Sample 81 (1.05e-4)
Ours vs. VAE 68 (3.28e-2)
Ours vs. Ground Truth 47 (7.11e-1)
Random vs. Ground Truth 34 (9.77e-3)
TABLE IV: User study results. The percentage of answers where Method A is preferred over Method B is listed. The numbers in the brackets denote p-value.
@@ V-C2 Perceptual User Study
For more comprehensive evaluation, we have conducted a perceptual user study to evaluate the realism of the animated 3D eye gaze. We employed an A/B test, presenting comparisons in a random order to evaluate our TalkingEyes method against four baseline methods and ground truth recordings. In each comparison, the participants were presented with two video clips with audio and asked: “Which recording exhibits more natural eyeball movements with fewer artifacts, such as speech-unrelated motions, exaggerated movements, or static poses?” They responded by selecting either A or B. For the study, we randomly selected 20 audio clips from TKED-Test, and used them to generate animations for six kinds of A/B comparisons (5 audios $\times$ 6 comparisons). 20 participants took part in the study, finally yielding 600 entries. We calculated the preference rate for Method A by determining the percentage of times it was chosen over Method B in a pairwise comparison. The results of this A/B testing are tabulated in Tab. IV. We also conducted Wilcoxon test in the study, in which the p-value less than 0.05 indicates a statistically significant difference between Method A and B.
As clearly shown in Tab. IV, participants preferred our method over all the four baselines. However, they still favored the ground truth recording over our method, which validates our experimental setup. The last comparison (“Random vs. Ground Truth”) confirms that participants can perceive the correspondence between speech and eye gaze motion, as the “Random” baseline pairs ground truth eye gaze motion with unrelated audio.
Figure 6: Visual comparisons of sampled frames of 3D facial animations generated by FaceFormer [FaceFormer2022], CodeTalker [CodeTalker2023], Learn2Talk [Learn2Talk2025] and our method.